A vessel is full of mixture of spirirt and water in which there is 18% spirit. 8 litres are drawn off and the vessel is filled up with water. If the spirit is now 15%, find the quantity in the vessel.
Answer by mananth(16946)

A vessel is full of mixture of spirirt and water in which there is 18% spirit. 8 litres are drawn off and the vessel is filled up with water. If the spirit is now 15%, find the quantity in the vessel.
..
let their be x litres in the vessel 18%
8 litres removed.
balance = x-8 --- 18%
water = 8 litres 0%
total x litres 15%
..
0.18(x-8)+0 = 0.15x
0.18x-1.44=0.15x
0.18x-0.15x=1.44
0.03x=1.44
x=48 litres